Math 2890 QZ-6 SP 2018 1) Find the rank of the following matrix. Also find a basis for the row and column spaces. 1 0 3 3 10 0 -1 2 Find a basis of Null(A) where A is the given matrix. Find the rank of A and dimension of Nul(A). Let B be an invertible 4X4 matrix (a matrix with 4 rows and 4 columns). Is the matrix AATB also invertible? Explain.
Answer #1

the given matrix is   1 0 3C0 3 1 10 0 1 12 1 1 1 2-2

reduced the matrix to row echlon form

R2\rightarrowR2-3R1 , R3\rightarrowR3+R1 ,R4\rightarrowR4-R1 we have

001 311 011 1000

apply R3\rightarrowR3-R2 ,R4\rightarrowR4+R2

2 001 3100 0100 1000

R4\rightarrowR4+2R3

0010 3100 0100 1000

which is the row echlon form having one complete zerro row

hence its nullity=1 and rank=4-1=3

## the matrix AATB is not invertible if B is invertible

{ a matrix is invertible if its determinant is not zero}

as rank of A is 3<4=order of the matrix

hence determinant of A=0

so det(AATB)=det(A)det(ATB)

=0.det(ATB)=0

hence it is not invertible